OpenStudy (anonymous):

vehicle A averages 14 miles per gallon of gasoline and Vehicle B averages 36 miles per gallon of gasoline. At these rates, how many more gallons of gasoline does Vehicle A need than Vehicle B to make a 1,008 mile trip

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

Vehicle A: 1008/ 14 = about 72 gallons Vehicle B: 1008/ 36 = about 28 gallons 72 - 28 = 40 ANSWER: about 44 gallons

OpenStudy (anonymous):

Sorry when I subtracted the 28 from 72 I meant 44, not 40.

OpenStudy (anonymous):

ah i see thank you

OpenStudy (anonymous):

You're Welcome!:)